OpenStudy (anonymous):

How do I determine the equation of the inverse of this function: f(x) = (x-4)^2+3

OpenStudy (anonymous):

@TuringTest, @Chlorophyll do you guys know how to do this? :$

OpenStudy (turingtest):

let y=f(x) so we get y=(x-4)^2+3 solve this for x and what do you get?

OpenStudy (anonymous):

so this is what I did: x= (y-4)^2+3 x-3 = (y-4)^2 x-3 = \[sqrt{y-4}\]

OpenStudy (anonymous):

and then I'm stuck..

OpenStudy (turingtest):

\[x-3=(y-4)^2\]square root of both sides gives\[\sqrt{x-3}=y-4\]

OpenStudy (anonymous):

ohhh okay shoot I forgot, thanks!

OpenStudy (turingtest):

welcome :)